Machinery purchased for $50,000 by Tom Brady Co. in 2010 was originally estimated to have a life of 10 years with a salvage value of $5,000 at the end of that time. Depreciation has been entered for 6 years on this basis. In 2016, it is determined that the total estimated life should be 12 years with a salvage value of $2,500 at the end of that time. Assume straight-line depreciation. 


Instructions 

(a) Determine the depreciation expense for 2016.
Answer #1

Answer

  • Requirement [a]
    Depreciation expense for 2016 = $ 3,417 (or $ 3416.67 or $ 3416.7)

A

Cost

$               50,000.00

B

Residual Value

$                 5,000.00

C=A - B

Depreciable base

$               45,000.00

D

Life [in years]

10

E=C/D

Annual SLM depreciation

$                 4,500.00

F = E x 6

6 year Accumulated Depreciation

$               27,000.00

G = A - F

Book value at the time of revision

$               23,000.00

H

New Residual Value

$                 2,500.00

I = G - H

New Depreciable base

$               20,500.00

J = 12 years - 6 years passed

New remaining life

6

K = I/J

New revised depreciation from 7th year

$ 3,416.67 Answer
Answer #2

(a) Annual depreciation = (50000-5000)/10 = $4500 per year

Net book value at the beginning of 2016 = $50000 - 4500*6 = $23,000

Depreciation expense for 2016 = (23000-2500)/6 = $3417